sum of 800 is in the denominatiom of rs 10 and rs 20 if total no of notes be 50 fund the no of noets of each type
Answer by Maths68(1474)

Let
Total Rs. 10 = x notes
Total Rs. 20 =50-x
Total Number of Notes = 50
Total sum of all notes = 800
(10)(x)+(50-x)(20)=800
10x+1000-20x=800
-10x=800-1000
-10x=-200
Divide by -10
-10x/-10=-200/10
x=20

Total Rs. 10 = x = 20 notes
Total Rs. 20 =50-x = 50-20 = 30 notes